The stock of XYZ sells for $65 a share. Its likely dividend payout and end-of-year price depend on the state of the economy by the end of the year as follows:

Dividend Stock Price
Boom $2.40 $73
Normal economy 1.60 66
Recession 0.85 57

a. Calculate the expected holding-period return and standard deviation of the holding-period return. All three scenarios are equally likely. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

b. Calculate the expected return and standard deviation of a portfolio invested half in XYZ stock and half in Treasury bills. The return on bills is 3%. (Do not round intermediate calculations. Round your answers to 2 decimal places.)
Answer #1

1.
=1/3*((2.40+73)/65-1)+1/3*((1.60+66)/65-1)+1/3*((0.85+57)/65-1)
=3.0000%

2.
=sqrt(1/3*((2.40+73)/65-1-3.0000%)^2+1/3*((1.60+66)/65-1-3.0000%)^2+1/3*((0.85+57)/65-1-3.0000%)^2)
=11.0454%

3.
=3.0000%

4.
=1/2*11.0454%
=5.5227%